OpenStudy (anonymous):

The salad dressing comes in two different size bottles. The 16 ounce bottle is on sale for $3.36. The 10 ounce bottle is priced at $2.30. Darius has a coupon for 40 cents off if you buy two 10 ounce bottles. Determine if Darius gets the better buy when he buys one 16 ounce bottle or when he buys two 10 ounce bottles. a. Darius will get the same deal on either purchase. b. Darius will get the better deal buying one 16 ounce bottle. c. Darius will get the better deal buying two 10 ounce bottles. d. Darius doesn’t need the coupon, he will get the best deal on one 10 ounce bottle

OpenStudy (anonymous):

Okay, so 1 bottle of 10 oz is 2.30*2=4.60 - the 40 cent coupon, is 4.20, which is more expensive than the 1 16 oz bottle. so the answer is b. I'm pretty sure.. If I'm wrong I apologize.

OpenStudy (jdoe0001):

so if he buys the 16oz bottle, he'd pay $3.36 if he buys the 2 10oz bottles, he'd pay $2.30 * 2 - 4(from the coupon) so which of the two gives the cheaper value, that is, less dollar per ounce, well \(\large \begin{array}{rllll} \cfrac{3.36dollars}{16oz}\implies \square \frac{dollars}{oz}\quad ?\\ \quad \\ \cfrac{(2.3\cdot 2)-4dollars}{20oz}\implies \square \frac{dollars}{oz}\quad ? \end{array}\) so... which one gives the cheaper ounce per dollar?